OpenStudy (anonymous):

approximate (1.999)^4 using differentials

OpenStudy (tkhunny):

\(y = x^{4}\) \(dy = 4x^{3}dx\) \(2^{4} = 16\) \(2 - 0.001 = 1.999\) \(y + dy = x^{4} + 4x^{3}dx\) \(y + dy = 16 + 4(8)(-0.001)\) Make any sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (tkhunny):

Sweet! Now, you do one.